Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

At a blind taste competition a contestant is offered 3 cups

Expert replies
by varun289 » Sat Apr 13, 2013 10:48 am
At a blind taste competition a contestant is offered 3 cups of each of the 3 samples of tea in a random arrangement of 9 marked cups. If each contestant tastes 4 different cups of tea, what is the probability that a contestant does not taste all of the samples?

A. \frac{1}{12}

B. \frac{5}{14}

C. \frac{4}{9}

D. \frac{1}{2}

E. \frac{2}{3}
Join the discussion
Source: — Problem Solving |

by Anju@Gurome » Sat Apr 13, 2013 11:21 am
varun289 wrote:At a blind taste competition a contestant is offered 3 cups of each of the 3 samples of tea in a random arrangement of 9 marked cups. If each contestant tastes 4 different cups of tea, what is the probability that a contestant does not taste all of the samples?
Probability that a contestant does not taste all of the samples = 1 - probability that a contestant tastes all 3 samples.

This means first we should find the probability that a contestant tastes 2 cups of one sample and 1 cup from each of other 2 samples.

Total number of ways to choose 4 cups out of 9 = 9C4 = 9!/[(4!)*(5!)] = 9*8*7*6/4! = 9*2*7
Number of ways to choose the sample which provides with 2 cups = 3C1 = 3
Number of ways to choose these 2 cups from the sample chosen = 3C2 = 3
Number of ways to choose 1 cup out of 3 from the 2nd sample = 3C1 = 3
Number of ways to choose 1 cup out of 3 from the 3rd sample = 3C1 = 3

So, probability that a contestant tastes all 3 samples = (3*3*3*3)/(9*2*7) = 9/14
Therefore, required probability = 1 - 9/14 = 5/14

The correct answer is B.
Anju Agarwal
Quant Expert, Gurome

Backup Methods : General guide on plugging, estimation etc.
Wavy Curve Method : Solving complex inequalities in a matter of seconds.

§ GMAT with Gurome § Admissions with Gurome § Career Advising with Gurome §
Join the discussion

by J N » Fri Apr 19, 2013 3:26 pm
how would/or is it possible to do this with 1-prob of getting all 3

choose A, then B, then C then any

3/9 * 3/8 * 3/7 * 6/6

take this and multiple by 3! to get different abc arrangements

it does not work but wondering if can do it and how
Join the discussion

by GMATGuruNY » Sat Apr 20, 2013 4:18 am
At a blind taste competition a contestant is offered 3 cups of each of the 3 samples of tea in a random arrangement of 9 marked cups. If each contestant tastes 4 different cups of tea, what is the probability that a contestant does not taste all of the samples?

A. 1/12

B. 5/14

C. 4/9

D. 1/2

E. 2/3

P(NOT selecting all 3 types of tea) = 1 - P(selecting all 3 types of tea).


Let the 3 types of tea be A, B and C.
For all 3 types of tea to be tasted, one type must be selected EXACTLY TWICE.

P(exactly n times) = P(one way) * (total possible ways).

One way to select A exactly twice: AABC
P(A is selected 1st) = 3/9. (Of the 9 cups, 3 contain A.)
P(A is selected 2nd) = 2/8. (of the 8 remaining cups, 2 contain A.)
P(B is selected 3rd) = 3/7. (Of the 7 remaining cups, 3 contain B.)
P(C is selected 4th) = 3/6. (Of the 6 remaining cups, 3 contain C.)
Since we want all of these events to happen, we multiply the probabilities:
3/9 * 2/8 * 3/7 * 3/6 = 1/56.

Total possible ways to select A exactly twice:
AABC is only ONE WAY for A to be selected exactly twice.
Since arrangement of AABC represents a different way to select A exactly twice, the result above must be multiplied by the number of ways to arrange AABC.
The number of ways to arrange 4 elements = 4!.
But when an arrangement includes IDENTICAL elements, we must divide by the number of ways to arrange the identical elements.
The reason: when the identical elements swap positions, the arrangement doesn't change.
Since AABC includes two identical elements, we divide by 2!:
Number of ways to arrange AABC = 4!/2! = 12.
Thus:
Number of ways to select A exactly twice = 1/56 * 12 = 3/14.

BBAC and CCAB:
Since any of the 3 teas could be the one selected exactly twice, the result above must be multiplied by 3:
Thus:
P(selecting all 3 types of tea) = 3/14 * 3 = 9/14.

Thus:
P(not selecting all 3 types of tea) = 1 - 9/14 = 5/14.

The correct answer is B.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by Amrabdelnaby » Fri Nov 27, 2015 10:12 am
Hi Guru,

can you please show me how to solve it using combinations?

Thanks
GMATGuruNY wrote:
At a blind taste competition a contestant is offered 3 cups of each of the 3 samples of tea in a random arrangement of 9 marked cups. If each contestant tastes 4 different cups of tea, what is the probability that a contestant does not taste all of the samples?

A. 1/12

B. 5/14

C. 4/9

D. 1/2

E. 2/3

P(NOT selecting all 3 types of tea) = 1 - P(selecting all 3 types of tea).


Let the 3 types of tea be A, B and C.
For all 3 types of tea to be tasted, one type must be selected EXACTLY TWICE.

P(exactly n times) = P(one way) * (total possible ways).

One way to select A exactly twice: AABC
P(A is selected 1st) = 3/9. (Of the 9 cups, 3 contain A.)
P(A is selected 2nd) = 2/8. (of the 8 remaining cups, 2 contain A.)
P(B is selected 3rd) = 3/7. (Of the 7 remaining cups, 3 contain B.)
P(C is selected 4th) = 3/6. (Of the 6 remaining cups, 3 contain C.)
Since we want all of these events to happen, we multiply the probabilities:
3/9 * 2/8 * 3/7 * 3/6 = 1/56.

Total possible ways to select A exactly twice:
AABC is only ONE WAY for A to be selected exactly twice.
Since arrangement of AABC represents a different way to select A exactly twice, the result above must be multiplied by the number of ways to arrange AABC.
The number of ways to arrange 4 elements = 4!.
But when an arrangement includes IDENTICAL elements, we must divide by the number of ways to arrange the identical elements.
The reason: when the identical elements swap positions, the arrangement doesn't change.
Since AABC includes two identical elements, we divide by 2!:
Number of ways to arrange AABC = 4!/2! = 12.
Thus:
Number of ways to select A exactly twice = 1/56 * 12 = 3/14.

BBAC and CCAB:
Since any of the 3 teas could be the one selected exactly twice, the result above must be multiplied by 3:
Thus:
P(selecting all 3 types of tea) = 3/14 * 3 = 9/14.

Thus:
P(not selecting all 3 types of tea) = 1 - 9/14 = 5/14.

The correct answer is B.
Join the discussion

by Brent@GMATPrepNow » Fri Nov 27, 2015 10:33 am
At a blind taste competition a contestant is offered 3 cups of each of the 3 samples of tea in a random arrangement of 9 marked cups. If each contestant tastes 4 different cups of tea, what is the probability that a contestant does not taste all of the samples?

A. 1/12
B. 5/14
C. 4/9
D. 1/2
E. 2/3
This question can be solved using the complement.
That is, P(Event A happening) = 1 - P(Event A not happening)
So, here we get: P(contestant does not taste all 3 samples) = 1 - P(contestant DOES taste all 3 samples)

P(contestant DOES taste all 3 samples)
For this event to occur, the contestant must taste 2 cups of one sample, 1 cup from another sample, and 1 cup from another sample.

Let's take the task of tasting all 3 samples and break it into STAGES.
Stage 1: Select the tea that will be tasted twice. There are 3 types of tea, so stage 1 can be completed in 3 ways.

Stage 2: Choose 2 cups to taste from tea selected in stage 1. Since the order in which we select the 2 cups does not matter, we can use combinations. We can select 2 cups from 3 cups in 3C2 ways(= 3 ways).

Stage 3: From one of the two remaining (untasted) teas, select 1 cup to taste. There are 3 cups, so stage 3 can be completed in 3 ways.

Stage 4: Select 1 cup from the last remaining (untasted) tea. There are 3 cups, so stage 4 can be completed in 3 ways.

By the Fundamental Counting Principle (FCP), we can complete all 4 stages in (3)(3)(3)(3) ways (= 81 ways)

The TOTAL number of ways to select 4 cups from 9 cups = 9C4 = 126

So, P(contestant DOES taste all 3 samples) = 81/126 = 9/14

This means that P(contestant does not taste all 3 samples) = 1 - 9/14
= [spoiler]5/14 [/spoiler]
= B

Note: the FCP can be used to solve the MAJORITY of counting questions on the GMAT. For more information about the FCP, watch our free video: https://www.gmatprepnow.com/module/gmat- ... /video/775

Then you can try solving the following questions:

EASY
- https://www.beatthegmat.com/what-should- ... 67256.html
- https://www.beatthegmat.com/counting-pro ... 44302.html
- https://www.beatthegmat.com/picking-a-5- ... 73110.html
- https://www.beatthegmat.com/permutation- ... 57412.html
- https://www.beatthegmat.com/simple-one-t270061.html
- https://www.beatthegmat.com/mouse-pellets-t274303.html


MEDIUM
- https://www.beatthegmat.com/combinatoric ... 73194.html
- https://www.beatthegmat.com/arabian-hors ... 50703.html
- https://www.beatthegmat.com/sub-sets-pro ... 73337.html
- https://www.beatthegmat.com/combinatoric ... 73180.html
- https://www.beatthegmat.com/digits-numbers-t270127.html
- https://www.beatthegmat.com/doubt-on-sep ... 71047.html
- https://www.beatthegmat.com/combinatoric ... 67079.html


DIFFICULT
- https://www.beatthegmat.com/wonderful-p- ... 71001.html
- https://www.beatthegmat.com/ps-counting-t273659.html
- https://www.beatthegmat.com/permutation- ... 73915.html
- https://www.beatthegmat.com/please-solve ... 71499.html
- https://www.beatthegmat.com/no-two-ladie ... 75661.html
- https://www.beatthegmat.com/laniera-s-co ... 15764.html

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by Brent@GMATPrepNow » Fri Nov 27, 2015 10:36 am
At a blind taste competition a contestant is offered 3 cups of each of the 3 samples of tea in a random arrangement of 9 marked cups. If each contestant tastes 4 different cups of tea, what is the probability that a contestant does not taste all of the samples?

a) 1/12
b) 5/14
c) 4/9
d) 1/2
e) 2/3
Let's say that there are 3 kinds of tea: A, B and C, and there are 3 cups of each tea.

First find P(contestant does not taste tea A)
P(contestant does not taste tea A) = P(1st selection is not tea A AND 2st selection is not tea A AND 3rd selection is not tea A AND 4th selection is not tea A)
= P(1st selection is not tea A) x P(2st selection is not tea A) x P(3rd selection is not tea A) x P(4th selection is not tea A)
= 6/9 x 5/8 x 4/7 x 3/6
= 5/42

Now find P(contestant does not taste tea B)
The steps to find this probability will be the same as steps taken to find the above probability.
So, P(contestant does not taste tea B) = 5/42

Likewise, P(contestant does not taste tea C) = 5/42

So, P(contestant does not taste all of the samples) = 5/42 + 5/42 + 5/42
=[spoiler] 5/14 [/spoiler]
= B

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion